# Chain-of-Influence: Tracing Interdependencies Across Time and Features in Clinical Predictive Modelings

### v1

@@ 3.3 Cross-Feature Attention and Chained Influence
To integrate the local feature importance with temporal relationships, we define a chained influence measure that quantifies the effect of an earlier feature on a later one:
$\text{I}(t,i;t^{\prime},j)=C[t,i]\times A[t,t^{\prime}]\times C[t^{\prime},j],$
where $C[t,i]$ represents the local contribution of feature $i$ at time $t$, $C[t^{\prime},j]$ represents the local contribution of feature $j$ at time $t^{\prime}$, and $A[t,t^{\prime}]$ denotes the attention weight that quantifies the extent to which information from time $t$ is incorporated into the representation at time $t^{\prime}$. This formulation effectively captures the dynamic interplay between individual feature contributions and their propagation through time, offering a detailed and interpretable account of how early features influence later ones within the CoI framework.
3.4 Implementation Optimizations
To achieve high computational efficiency and robust performance, we introduce two key implementation optimizations:
Vectorized Computation of Temporal-Feature Cross Influence: To compute the chained influence tensor efficiently, we leverage array stacking and the Einstein summation convention via np.einsum. This vectorized approach replaces nested loops with concise batch-level operations, dramatically improving computational efficiency and scalability. Specifically, the influence tensor is computed as:
$\mathrm{I}[n,t,u,i,j]=A[n,t^{\prime},t]\times C[n,t,i]\times C[n,t^{\prime},j],$
where $n$ indexes samples, $t$ and $t^{\prime}$ index time steps (with $t<t^{\prime}$), and $i$ and $j$ index features.
Dynamic Tanh (DyT) Integration: We replace traditional normalization layers with Dynamic Tanh (DyT) activation function Zhu et al. (2025), which dynamically adapts the non-linearity based on the input distribution:
$\text{DyT}(x)=\alpha\cdot\tanh(\beta\cdot x+\gamma),$
where $\alpha$, $\beta$, and $\gamma$ are learnable parameters.
4 Experimental Setup
4.1 Datasets
We evaluate our Chain-of-Influence framework on two complementary clinical datasets representing distinct prediction scenarios: chronic disease progression and acute care outcomes.
@@ 4.1.1 Chronic Kidney Disease (CKD) Dataset
Our primary evaluation utilizes a proprietary longitudinal CKD dataset capturing gradual kidney function deterioration over extended periods, making it ideal for evaluating temporal-feature interdependencies in chronic disease progression.
The cohort comprises 1,422 CKD patients monitored over 24 months at 3-month intervals (8 time points), with 38 clinical features spanning demographics, comorbidities, laboratory biomarkers, and healthcare utilization. The prediction target is progression to End-Stage Renal Disease (ESRD), occurring in 6.0% of patients—a critical endpoint requiring dialysis or transplantation.
@@ 4.1.2 MIMIC-IV Dataset
To demonstrate generalizability, we evaluate on MIMIC-IV v3.1 Johnson et al. (2023), a public critical care database. From 94,458 ICU stays, we constructed a cohort of 65,366 patients using their first ICU stay, extracting hourly measurements over 48 hours across 15 clinical features (vital signs, laboratory tests, demographics).
Unlike the CKD dataset’s monthly resolution, MIMIC-IV provides hourly temporal data capturing rapid physiological changes characteristic of critical illness. The prediction target is in-hospital mortality (10.8% prevalence).
These complementary datasets enable evaluation across different temporal scales (hours vs. months) and clinical contexts (acute vs. chronic care). Comprehensive dataset characteristics, feature specifications, and preprocessing details are provided in Appendix A.
@@ 4.2 Baselines
We compare Chain-of-Influence (CoI) against three representative baselines to benchmark its performance. We include a standard BiLSTM as a strong but non-interpretable recurrent baseline that processes sequences bidirectionally. To compare against models with explicit temporal attention, we use RETAIN, the canonical interpretable baseline with a dual-level attention mechanism. Finally, to assess CoI against a different attention paradigm with strong clinical priors, we use StageNet, which leverages channel-wise attention to re-weight features based on disease stage. This selection provides a comprehensive benchmark against a standard recurrent model (BiLSTM), a model with explicit temporal attention (RETAIN), and one focused on channel-wise attention (StageNet).
@@ 4.3 Evaluation Metrics
We evaluate model performance using standard classification metrics. The primary metric is the Area Under the Receiver Operating Characteristic curve (AUROC) Hanley & McNeil (1982); Fawcett (2006), which measures the model’s discriminative ability across all classification thresholds. We also report the F1-Score Van Rijsbergen (1979), the harmonic mean of precision and recall that provides balanced performance assessment for imbalanced datasets. Additionally, we assess overall prediction accuracy Sokolova & Lapalme (2009), which captures the proportion of correctly classified instances across both classes. Finally, we examine precision Van Rijsbergen (1979), which measures positive predictive value and is particularly important in healthcare applications to minimize false positive predictions Powers (2011).
@@ 5.1 Predictive Performance
Table 1: Performance comparison across datasets.
The comparative evaluation demonstrates the robust performance of Col across two distinct clinical prediction tasks. On the chronic CKD dataset, Col establishes clear superiority, outperforming the next-best model, StageNet, across all metrics. It achieves a notable 5.3% improvement in F1-score ( 0.721 vs. 0.685 ) and a $5.1\%$ increase in precision ( 0.820 vs. 0.780 ). This highlights Col’s enhanced ability to model the slow, interconnected progression of chronic disease over long timeframes.
On the acute care MIMIC-IV dataset, Col again delivers state-of-the-art performance, leading in AUROC, F1-score, Accuracy, and Precision. The results present a clinically significant nuance when comparing Col to RETAIN. While RETAIN achieves perfect recall (1.000), it does so with very low precision (0.500), indicating a high rate of false positives. In a critical care setting, such a high false alarm rate can lead to unnecessary interventions and alarm fatigue. In contrast, Col achieves a much stronger and more balanced profile, with high precision ( 0.850 ) and recall (0.880), culminating in a far superior F1-score (0.865 vs. RETAIN’s 0.667).
The performance of the baselines reveals key architectural insights. The standard Bi-LSTM performs adequately on the CKD data but struggles significantly with the complex, high-frequency interactions in the MIMIC-IV dataset ( $\mathrm{F}1=0.267$ ), showcasing the limitations of standard recurrent models in acute care. RETAIN’s behavior suggests its dual-attention mechanism is highly sensitive in the acute setting, adept at catching all true cases but at the cost of clinical utility due to low precision. StageNet proves to be a strong and balanced baseline on both datasets, making Col’s performance gains over it particularly meaningful.
Ultimately, Col’s strength lies in its architectural robustness. Its ability to achieve top-tier performance in both a sparse, long-term chronic setting and a dense, short-term acute care setting demonstrates that explicitly modeling feature influence chains captures generalizable patterns of deterioration. This suggests our framework has broad applicability across diverse clinical domains.
@@ 5.2 Interpretability Analysis
Due to space constraints, we present Chain-of-Influence’s interpretability capabilities using the CKD dataset, whose complex temporal dynamics and extended observation periods provide an ideal framework for demonstrating influence chain analysis.
5.2.1 Temporal Attention Patterns
Figure 1: Temporal attention comparison between RETAIN and CoI models. CoI (red) shows pronounced attention focusing on final time steps ($t_{6}$-$t_{7}$) with peak weights reaching 0.29, while RETAIN (blue) maintains more uniform distribution with maximum attention of 0.22 at $t_{7}$. The U-shaped pattern of CoI reflects clinical understanding that both baseline measurements and pre-progression periods are most informative for ESRD prediction.
Figure 1 presents a striking comparison of temporal attention patterns between CoI and RETAIN across the 8-time-point CKD progression timeline. The visualization reveals fundamentally different attention strategies: while RETAIN exhibits relatively uniform attention distribution with gradual increases toward recent time points, CoI demonstrates dramatic temporal focusing with a distinct U-shaped pattern that culminates in peak attention at the final time steps.
The temporal attention analysis reveals several clinically significant patterns. Most notably, CoI exhibits minimal attention weights (approximately 0.08) during the middle monitoring periods ($t_{1}$-$t_{3}$), followed by a dramatic escalation beginning at $t_{4}$ and reaching maximum attention at $t_{6}$-$t_{7}$ (0.29). This attention pattern aligns remarkably well with clinical understanding of CKD progression, where the final monitoring periods before ESRD onset contain the most prognostically valuable information. In contrast, RETAIN’s attention increases more gradually from 0.09 at $t_{1}$ to 0.22 at $t_{7}$, failing to distinguish between routine monitoring periods and critical pre-progression phases.
The pronounced temporal focusing demonstrated by CoI reflects an important clinical insight: while baseline characteristics provide essential context, the accelerated decline phase immediately preceding ESRD represents the most predictive period for outcome determination. This pattern suggests that CoI successfully learns to identify the temporal windows when cascading physiological changes become irreversible, enabling more precise risk stratification than models with uniform temporal weighting.
5.2.2 Feature Importance Analysis
(a) Feature importance scores for RETAIN model on CKD dataset.
(b) Feature importance scores for CoI model on CKD dataset.
Figure 2: Comparison of feature attributions on CKD. Normalized feature-importance profiles for (a) RETAIN and (b) CoI on the CKD dataset, shown side-by-side to compare salient predictors and attribution sparsity. Values are on the same scale to enable direct visual comparison; higher bars indicate greater contribution to the model’s decision.
While traditional feature importance rankings provide valuable insights, our Chain-of-Influence approach reveals the limitations of static importance scores in capturing temporal dependencies. Figure 2 presents detailed feature importance comparisons between RETAIN and CoI, showing both models consistently identify diabetes as the primary risk factor, corroborating established clinical evidence Burckhardt et al. (2017); Li et al. (2025a). Laboratory markers (S4, S5), healthcare expenditures (net_exp_P/O), and hemoglobin levels consistently rank among top features across both models.
However, static importance rankings obscure the temporal dynamics that CoI’s influence chains reveal. While CoI appropriately ranks eGFR higher ($6^{th}$) than RETAIN ($14^{th}$), neither model captures how eGFR’s influence propagates through temporal cascades to drive healthcare utilization patterns. Similarly, traditional rankings show diabetes as most important but fail to reveal how its influence chains through cardiovascular complications to accelerate kidney function decline over specific time horizons.
The Chain-of-Influence paradigm transcends these limitations by mapping how feature importance evolves temporally. Rather than asking which features are most important, CoI reveals when features become influential and how their effects cascade through clinical pathways. This temporal perspective explains why features like age demonstrate broad systemic influence across multiple chains while race shows targeted effects within specific progression pathways, providing clinicians with actionable insights beyond static importance scores.
5.2.3 Chain-of-Influence Analysis
Figure 3 reveals CoI’s breakthrough capability to map chains of influence across the 24-month CKD progression timeline. The network exposes how clinical deterioration propagates through interconnected pathways: cardiovascular complications (CHF_t-6, CD_t-5) trigger cascading effects through kidney function decline (eGFR_t-5, eGFR_t-6) to advanced disease stages (S4_t-5, S5_t-6), ultimately culminating in healthcare system escalation (n_claims_I_t-4, net_exp_DR_t-6). This temporal chaining quantifies the established cardio-renal syndrome with unprecedented precision, transforming clinical intuition into measurable influence pathways.
Critically, CoI identifies persistent influence chains that capture the irreversible nature of advanced clinical conditions. The S5_t-5 → S5_t-6 → S5_t-7 progression exemplifies how CKD Stage 5 creates self-reinforcing temporal chains, where each occurrence dramatically amplifies subsequent disease burden and progression risk. Similarly, CHF demonstrates persistent chaining patterns (CHF_t-6 → CHF_t-4 → CHF_t-1), reflecting the chronic, progressive nature of heart failure that compounds kidney disease severity. These persistent chains transcend simple temporal adjacency—they represent traced pathways of irreversible pathophysiological damage that fundamentally alter patient trajectories, enabling CoI to distinguish between transient clinical events and permanent disease state transitions.
The influence chains demonstrate remarkable bidirectional complexity, particularly evident in healthcare utilization patterns. CoI discovers that increased outpatient visits (n_claims_O_t-7) and professional consultations (n_claims_P_t-6) form feedback loops—simultaneously reflecting disease progression while driving improved outcomes through enhanced monitoring. This bidirectional influence challenges traditional unidirectional prediction models and reveals how clinical interventions reshape disease trajectories through temporal chains of care optimization.
CoI’s selective demographic integration further validates the chain-of-influence paradigm. Age propagates broad systemic effects across multiple clinical pathways, while gender and race demonstrate targeted, condition-specific influence patterns within discrete chains. The red-blue edge visualization immediately distinguishes risk-enhancing from protective influence chains, providing clinicians with the first quantitative map of CKD’s temporal complexity. This represents a paradigm shift from static risk scoring to dynamic influence tracking, enabling precise intervention timing based on evolving chain dynamics.
6 Limitations

### v2

@@ 3.2.3 Chain of Influence Measure
By integrating local contributions with temporal connectivity, we define a chained influence measure that quantifies how feature $i$ at time $t$ influences feature $j$ at time $t^{\prime}$ (with $t<t^{\prime}$):
$\mathcal{I}(t,i;t^{\prime},j)=C[t,i]\times A[t,t^{\prime}]\times C[t^{\prime},j].$ (15)
For efficient computation across all samples, time pairs, and feature pairs simultaneously, we implement this using vectorized Einstein summation.
4 Experiments
4.1 Data
We evaluate Chain-of-Influence on two complementary clinical settings that differ in both clinical tempo (chronic vs. acute) and temporal structure (regular vs. irregular sampling).
@@ 4.1.1 Chronic Kidney Disease (CKD) Dataset
Our primary evaluation uses a proprietary longitudinal CKD cohort with scheduled nephrology follow-ups. By protocol, patients revisit every three months, yielding a regularly sampled sequence of 1,422 patients over 24 months (8 visits), with 38 features spanning demographics, comorbidities, laboratory biomarkers, and utilization. The prediction target is progression to End-Stage Renal Disease (ESRD), with 6.0% prevalence.
@@ 4.1.2 MIMIC-IV Dataset
To assess generalizability in fast-changing conditions, we use MIMIC-IV v3.1 johnson2023mimic and construct a cohort from first ICU stays (65,366 patients out of 94,458 stays). Raw ICU measurements are recorded at irregular times; for modeling, we aggregate events into 1-hour windows over the first 48 hours and retain irregularity signals via masks and time-gap features (time since last observation) across 15 variables (vitals, labs, demographics). The outcome is in-hospital mortality (10.8% prevalence).
Together, these datasets stress-test CoI across chronic vs. acute trajectories and regular vs. irregular temporal structure: CKD offers fixed revisit schedules for long-horizon progression, whereas MIMIC-IV captures volatile ICU dynamics with uneven sampling. Detailed cohort definitions, feature specifications, and preprocessing steps are provided in the Appendix A.
@@ 4.2 Baselines
We benchmark Chain-of-Influence (Col) against four representative models: (i) BiLSTM, a strong but noninterpretable recurrent baseline that processes sequences bidirectionally; (ii) RETAIN, the canonical dual-attention model with visit-level ( $\alpha_{t}$ ) and feature-level ( $\beta_{t}$ ) attention; (iii) StageNet, which applies stageaware, channel-wise reweighting tailored to clinical progression; and (iv) BEHRT-style Transformer: since BEHRT was designed for tokenized medical-code inputs, we adapt its architecture backbone to our longitudinal, visit-level continuous features input settings.
@@ 4.3 Evaluation Metrics
We evaluate model performance using standard classification metrics. The primary metric is the Area Under the Receiver Operating Characteristic curve (AUROC) hanley1982meaning; fawcett2006introduction, which measures the model’s discriminative ability across all classification thresholds. We also report the F1-Score van1979information, the harmonic mean of precision and recall that provides balanced performance assessment for imbalanced datasets. Additionally, we assess overall prediction accuracy sokolova2009systematic, which captures the proportion of correctly classified instances across both classes. Finally, we examine precision van1979information, which measures positive predictive value and is particularly important in healthcare applications to minimize false positive predictions powers2011evaluation.
4.4 Training and Implementation Details
Before training, we impute missing values using a MICE-based technique raghunathan2001multivariate; van2007multiple applied separately on each dataset. Given the substantial class imbalance in both cohorts, we evaluate several mitigation strategies—including SMOTE, ADASYN, class-weighted loss, and Temporal SMOTE (TSMOTE)—and find that TSMOTE offers the best F1–AUROC trade-off (Detailed result is provided in Appendix C.2), so we adopt it for all main experiments. Hyperparameters are selected via grid search over model-specific search spaces (Appendix B). All models are optimized with Adam and a numerically stable BCEWithLogitsLoss, with early stopping based on validation AUROC. We report test performance as the mean and 95% confidence interval over five random seeds. Additional training details are provided in Appendix B.6.
@@ 5.1 Predictive Performance
Table 1 presents the comparative evaluation of CoI against four representative baselines across both clinical prediction tasks. CoI achieves state-of-the-art performance across both datasets, demonstrating the effectiveness of explicitly modeling temporal-feature influence chains. On the chronic CKD dataset, CoI attains the highest AUROC (0.960) and F1-score (0.721), outperforming the strong BEHRT-style Transformer baseline by +0.5% AUROC and +0.8% F1. While these gains may appear modest, they represent meaningful improvements over an already competitive transformer-based architecture, achieved through CoI’s dual-attention mechanism that explicitly decomposes temporal and feature-level importance.
@@
Table 1: Performance comparison across datasets. Full results with confidence intervals are provided in Appendix D.1
On the acute care MIMIC-IV dataset, CoI establishes clearer superiority with AUROC of 0.950 and F1 of 0.865, improving upon BEHRT-style by +3.7% F1. A clinically critical pattern emerges when comparing CoI to RETAIN: while RETAIN achieves perfect recall (1.000), it does so at severe cost to precision (0.500), resulting in a 50% false-positive rate that would trigger unnecessary interventions and alarm fatigue in ICU settings. CoI maintains strong recall (0.880) while achieving substantially higher precision (0.850), yielding a balanced profile essential for clinical deployment. This represents a 30% F1 improvement over RETAIN (0.865 vs. 0.667) while avoiding the trap of sensitivity-only optimization. Baseline performance reveals key architectural insights. Standard Bi-LSTM struggles dramatically with MIMIC-IV’s complex temporal dynamics (F1=0.267), exposing fundamental limitations of recurrent architectures without explicit attention. StageNet provides competitive performance through stage-aware modeling but falls short of transformer-based approaches on both tasks. BEHRT-style Transformer emerges as the strongest baseline, validating self-attention’s effectiveness for clinical sequences—making CoI’s consistent improvements particularly significant. That CoI outperforms a strong transformer baseline suggests that its hierarchical attention—separating temporal selection, feature emphasis, and cross-feature interactions—captures richer progression patterns than vanilla self-attention, a hypothesis we examine more thoroughly via ablation study in Sec. 5.4.
@@ 5.2 Interpretability Analysis
5.2.1 Attention Pattern Comparison
Figure 2: Temporal and feature-level importance comparison. (a) Temporal attention comparison reveals CoI’s focused U-shaped pattern versus RETAIN’s uniform weighting. (b-c) Feature importance profiles demonstrate consistent identification of diabetes, laboratory markers (S4, S5), and healthcare utilization as primary predictors across both models.
Figure 3 summarizes CoI’s attention behavior in comparison to RETAIN on the CKD cohort. Figure 2(a) shows that both models learn a broadly U-shaped temporal importance pattern, placing higher weight on the baseline visit and on the final pre-ESRD visits. RETAIN, however, spreads mass relatively uniformly across the 8 visits, with only a mild recency bias (attention rising from roughly 0.09 at $t_{1}$ to 0.22 at $t_{7}$). CoI, in contrast, exhibits a much sharper U-shape: attention is low during the middle visits ($t_{1}$$t_{3}$), increases steeply from $t_{4}$, and peaks at the last two visits ($t_{6}$$t_{7}$, around 0.29). This pattern aligns with qualitative feedback from our collaborating nephrologists: both the baseline state and the accelerated pre-ESRD phase carry the strongest prognostic signal, whereas routine intermediate visits are less informative. CoI thus goes beyond generic recency bias and concentrates on clinically meaningful temporal windows where deterioration becomes effectively irreversible.
At the feature level (Figure 2(b)2(c)), both models consistently identify key predictors, with diabetes emerging as the dominant risk factor, in line with clinical evidence burckhardt2017multi; li2025towards. Laboratory markers (S4, S5), healthcare expenditures (net_exp_P/O), and hemoglobin levels also rank among the top features. However, these static rankings obscure the temporal dynamics that CoI’s influence chains reveal. For example, CoI ranks eGFR as 6th versus 14th in RETAIN, but neither ranking captures how eGFR’s influence propagates through time to drive utilization patterns, or how diabetes chains through cardiovascular complications to accelerate kidney decline over specific horizons. The Chain-of-Influence paradigm addresses this gap by mapping when features become influential and how their effects cascade through clinical pathways, providing actionable insights beyond static scores.
In summary, CoI and RETAIN show consistent temporal and feature-level trends, and these patterns align with feedback from our collaborating nephrologists and clinical data providers. To more thoroughly assess the faithfulness of these attributions, we next perform a deletion-based sensitivity analysis over temporal and feature dimensions.
5.2.2 Sensitivity Analysis of Temporal and Feature Contributions
[…]
Deletion strategy $k$ (visits) CKD MIMIC-IV
High-attention visits 1 0.42 $\pm$ 0.03 0.38 $\pm$ 0.04
Middle-attention visits 1 0.17 $\pm$ 0.05 0.15 $\pm$ 0.02
Low-attention visits 1 0.08 $\pm$ 0.01 0.09 $\pm$ 0.02
High-attention visits 3 0.89 $\pm$ 0.06 0.82 $\pm$ 0.05
Middle-attention visits 3 0.36 $\pm$ 0.06 0.32 $\pm$ 0.03
Low-attention visits 3 0.18 $\pm$ 0.04 0.16 $\pm$ 0.04
Temporal sensitivity.
[…]
Deletion strategy $k$ (features) CKD MIMIC-IV
High-attention features 1 0.51 $\pm$ 0.09 0.47 $\pm$ 0.07
Middle-attention features 1 0.21 $\pm$ 0.03 0.19 $\pm$ 0.09
Low-attention features 1 0.11 $\pm$ 0.06 0.10 $\pm$ 0.05
High-attention features 3 1.02 $\pm$ 0.03 0.96 $\pm$ 0.04
Middle-attention features 3 0.43 $\pm$ 0.04 0.38 $\pm$ 0.03
Low-attention features 3 0.22 $\pm$ 0.03 0.20 $\pm$ 0.03
High-attention features 5 1.43 $\pm$ 0.09 1.34 $\pm$ 0.08
Middle-attention features 5 0.63 $\pm$ 0.05 0.57 $\pm$ 0.04
Low-attention features 5 0.31 $\pm$ 0.03 0.28 $\pm$ 0.03
Feature sensitivity.
[…]
Taken together, these deletion experiments provide quantitative evidence that CoI’s temporal and feature-level attributions are strongly aligned with the model’s internal decision-making, even though they do not by themselves establish causal relations in the clinical sense.
5.3 Chain-of-Influence Analysis
Figure 3: Patient-specific Chain-of-Influence visualizations. (a) Top-33 influence graph for a single patient, showing the most influential feature–time pairs and their directed paths into the prediction node. (b) Path-centric view anchored at CKD Stage 5 at $t_{5}$, highlighting the high-influence chain through subsequent visits that drives the final ESRD risk. Additional visualization examples are provided in Appendix E.
Figure 3(a) visualizes the patient–specific Chain-of-Influence graph induced by Eq. equation 15 for a single CKD patient after the observation window. Each node corresponds to a feature–time pair $(i,t)$ and nodes are arranged from left (earlier visits) to right (later visits and the prediction node). Each edge encodes the magnitude of $\mathcal{I}(t,i;t^{\prime},j)$. Because the full graph is dense, we introduce a top-$k$ sparsification scheme for visualization: at the final visit $t_{7}$ we retain only the $k$ most influential incoming feature–time pairs; for each of these nodes we again retain only their top-$k$ predecessors at $t_{6}$, and so on recursively. Panel 3(a) shows the resulting sparse, time–unfolded influence network for $k=3$, $T_{start}=t_{4}$, revealing a small set of clinically influence pathways that contribute to the patient’s risk prediction.
To support path–centric reasoning, we also trace the forward influence of a single clinically meaningful starting point. In Figure 3(b), we anchor the visualization on CKD Stage 5 at visit $t_{5}$ and highlight only nodes and edges on high–influence paths emanating from this feature. Background nodes and edges are faded, while the active chain is emphasized in yellow (nodes) and orange (edges). This view reveals how advanced CKD at $t_{5}$ propagates through subsequent utilization and severity markers at $t_{6}$ and $t_{7}$, and ultimately into the prediction node. Clinicians can thus read off a concrete, patient–level narrative; for example, the Stage 5 $t_{5}\rightarrow t_{6}\rightarrow t_{7}$ progression exemplifies a self–reinforcing temporal chain in which each occurrence amplifies later disease burden and progression risk—consistent with the clinical reality that Stage 5 CKD is effectively irreversible and a dominant driver of ESRD outcome.
Together, these graph–centric (top-$k$) and path–centric views demonstrate how CoI turns the abstract influence tensor into structured, human–readable stories of full mapping leading to the predictions. Additional visualization examples are provided in Appendix E.
5.4 Ablation Study
[…]
Figure 4: CoI architecture overview. Detailed results are provided in Appendix D.2
Component synergy proves essential—the ”Temporal only” and ”Feature only” variants underperform the full model by 5-10% in F1-score, while the ”Transformer Only” baseline shows the worst degradation (CKD F1: -0.116, MIMIC-IV F1: -0.135). This demonstrates that self-attention alone cannot replicate the structured temporal-feature decomposition provided by our dual attention pathways. The full CoI model’s superior performance emerges from the multiplicative integration of attention mechanisms (Equations 710), enabling feature attention to guide what information enters the transformer while temporal attention weights the final representation. Replacing DyT with standard LayerNorm reduces performance uniformly (CKD: -2.6% F1, MIMIC-IV: -3.0% F1), validating DyT’s adaptive normalization for clinical data’s non-stationary distributions.
6 Conclusion
We introduced Chain-of-Influence (CoI), a deep learning framework that advances clinical predictive modeling by explicitly capturing and visualizing temporal-feature interdependencies. Through a hierarchical architecture integrating temporal attention, feature-level attention, and cross-feature transformer layers, CoI achieves state-of-the-art performance across heterogeneous clinical settings—attaining 0.960 AUROC on chronic kidney disease progression and 0.950 AUROC on ICU mortality prediction. Our ablation studies demonstrate that these gains emerge from the synergistic integration of attention mechanisms, with feature attention proving more critical for chronic disease and temporal attention dominating in acute care. Deletion-based sensitivity analyses further validate that CoI’s learned attributions faithfully reflect its internal decision process.
Beyond predictive accuracy, CoI transforms black-box predictions into structured, patient-specific influence chains that mirror clinical reasoning. By mapping not only what features matter but when they become influential and how their effects cascade through clinical pathways, CoI provides actionable insights that transcend static importance scores—offering clinicians a transparent and generalizable framework for understanding the complex interplay of clinical variables over time.
Limitations